OpenStudy (anonymous):

Help Please ? What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

HELP!!!!!!!!!

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

y-3=4(x+1) y=4x+7

OpenStudy (anonymous):

4x - y = -7 @sara17 ?

OpenStudy (anonymous):

yes that's right